OpenStudy (anonymous):

WILL GIVE MEDAL TO ANYONE WHO ANSWERS CORRECTLY! What is the solution of 6 over the quantity x plus 5. = 2

OpenStudy (anonymous):

6=2x+10

OpenStudy (anonymous):

incorrect. mutiple choice answers include: −8 −2 2 8

OpenStudy (anonymous):

its -2

OpenStudy (anonymous):

you said 6/(x+5)=2 right?

OpenStudy (anonymous):

6/-2=-3+5=2

OpenStudy (anonymous):

yep @davyjones

OpenStudy (anonymous):

its -2 plug it in the equation

OpenStudy (anonymous):

im going to check now

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

6/(x + 5) = 2/1 -- this is a proportion cross multiply 2(x + 5) = (6)(1) 2x + 10 = 6 2x = 6 - 10 2x = -4 x = -2